40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7852     Entry time: Tue Dec 18 16:37:17 2012
Author: Jamie 
Type: Update 
Category: Alignment 
Subject: Post vent, pre door removal alignment 

[Jenne, Manasa, Jamie]

Now that we're up to air we relocked the mode cleaner, tweaked up the alignment, and looked at the spot positions:

mcspot_post_vent.pdf

The measurements from yesterday were made before the input power was lowered.  It appears that things have not moved by that much, which is pretty good.

We turned on the PZT1 voltages and set them back to their nominal values as recorded before shut-down yesterday.  Jenne had centered IPPOS before shutdown (IPANG was unfortunately not coming out of the vacuum).  Now we're at the following value: (-0.63, 0.66).  We need to calibrate this to get a sense of how much motion this actually is, but this is not insignificant.

 

ELOG V3.1.3-